# Revisiting interpolating flows in $(1+1)$ hydrodynamics

### v1

@@ 1 Introduction
Relativistic hydrodynamics has been widely successful in describing collective phenomena in high energy QCD and early universe physics. Based on experimental data obtained from relativistic Heavy Ion Collisions (HICs) conducted both at the RHIC and LHC, we can confirm that a hot and dense nuclear matter is generated during the initial stages of these collisions, commonly referred to as Quark-Gluon Plasma (QGP). It is astonishing that many physical phenomena associated to QGP is well approximated by the almost perfect, strongly-coupled fluid picture of high energy particle processes in these regimes. The problem of solving hydrodynamics in $(1+1)$ dimensions has subsequently receieved particular attention since for very high energy particles, rotational invariance in the transverse directions make the flow effectively two dimensional. Depending on the rapidity dependence of such a flow, one could classify the associated heavy-ion collision processes.
In general, there are two broad classes of solutions deemed most important in this regime, namely the ones with and without manifest boost invariance. The boost invariant regime for a perfect fluid was developed by Bjorken [1] (see also [2]) in his seminal work, which applies beautifully to the central rapidity region of high energy collisions of relativistic heavy ions. Due to the symmetry protected nature, one can easily determine the four-velocity profile for the beam in question. Although elegant in formulation, the Bjorken boost invariant predictions are hardly ever important for experiments, where often Gaussian rapidity distributions dominate the data for the dense hot matter produced in high energy collisions, giving rise to what is know as the Landau flow [3]. This has been experimentally observed in high energy scattering events [4, 5, 6, 7, 8, 9] [^3] More extensive discussions on Landau flows appear in a number of works, including [10, 11, 12, 13, 14, 15, 16]. This list is by no means exhaustive and the interested reader is requested to look at the references of these.. However lucrative, exact analytic solutions of fluid equations in $(1+1)$ dimensions remain scarce, and general solutions spanning all parameter space is still hard to find even for benchmarking purposes.
@@ 2 The set-up: Hydrodynamic equations
With the light-cone coordinates the partial derivatives become, ±$\partial_{\pm}=\frac{\partial}{\partial z_{\pm}}\equiv\frac{1}{2}\left(\frac{\partial}{\partial t}\pm\frac{\partial}{\partial z}\right)$ and the hydrodynamic equation takes the form,
±$\partial_{\pm}T^{01}+\frac{1}{2}\partial_{+}(T^{11}\pm T^{00})-\frac{1}{2}\partial_{-}(T^{11}\mp T^{00})=0.$ (4)
@@
$y\equiv\frac{1}{2}\ln\left(\frac{\epsilon+p}{\epsilon-p}\right)=\frac{1}{2}\ln\left(\frac{1+v}{1-v}\right).$ (5)
Substituting the definition of velocity and (2) in (4) we get,
@@ 3 Interpolating flows: Boost and beyond
$y=\frac{1}{2}\ln\left(\frac{\epsilon+p}{\epsilon-p}\right)=\frac{1}{2}\ln\left(\frac{t+z}{t-z}\right).$ (9)
@@
[…]
@@
where, $l_{\pm}=\sqrt{A^{2}\ln{f_{\pm}}+1}$. Here, we have tried to keep the form of the expressions similar to the original work [17]; but, note that the definition of $l_{\pm}$ is different here, and so is the extra $\frac{g+1}{g}$ in the expression of pressure. But more importantly, these expressions can be seamlessly interpolated to the boost invariant regime. To round out this discussion, we show how the general solution based on (28) is different from the boost invariant regime, via plotting the value of $y/\eta$ for varying proper time and varying $A$ in figure (1) and (2).
@@
Figure 2: The ratio of $y/\eta$ has been plotted against the proper time $\tau$ for a constant $\eta$ and varying interpolation parameter $A$. Some we see that for smaller and smaller $A$ values the graph becomes flatter around 1 which is what we expect since we move towards the Bjorken flow limit.
@@ 4 Entropy flow
$\frac{dp}{d\epsilon}=\frac{s}{T}\frac{dT}{ds}=c_{s}^{2}\equiv\text{const.}$. (31)
@@
±$\partial_{\pm}\Phi(z_{+},z_{-})\equiv u^{\mp}T=T_{0}e^{-\theta\mp y}.$ (36)
@@
$=\frac{s_{0}}{\sqrt{2g}}\exp{\left[\frac{g-1}{2A^{2}}\left(\left(\frac{1}{g}-A^{2}\theta\right)+\left(\left(\frac{1}{g}+A^{2}\theta\right)^{2}-\frac{A^{4}y^{2}}{g}\right)^{1/2}-4\right)\right]}$
@@
Note that this is in general a gaussian Landau-like distribution in rapidity.
@@ 5 General harmonic flow solutions
As discussed in preceeding sections, the solution (26) and (28) are one parameter solutions for the equation (7). But we can actually do better. The anstaz (12) solves the equation (7) but a homogeneous version thereof,
@@
[…]
@@ 6 Discussions and conclusions
In this work, we discussed a refinement of the generalized fluid ansatz that unifies the Landau and Bjorken approaches to studying relativistic hydrodynamics. The full equations of motion can be solved for longitudinal flows effectively in $(1+1)$ dimensions, including free parameters which act as dials between different regimes of physical flows. We found out explicit expressions with arbitrary speed of sound for the physical observables in this case, and also worked out the entropy distribution in rapidity evaluated at different freeze-out temperatures. The harmonic property of the rapidity in hydrodynamic regime gives a large parameter space for viable solutions in this case.
This note was merely to clarify the construction of the interpolating ansatz between boost invariant and non-invariant regimes. There are various questions and intricacies that remain in the discussion of exact solutions in $(1+1)$ hydrodynamics. An immediate question, for example, is the nature of the free parameter $A$. In principle this parameter measures the distance between the space-time rapidity and the kinematical rapidity, which coincide in the case of boost invariance. A very interesting recent geometric perspective into the problem of Bjorken physics comes from Carrollian symmetries [22], which are basically speed of light going to zero limit of the relativistic cousin thereof [23, 24]. The highly ultrarelativistic nature of heavy-ion flows can be approximated as an effective $c\to 0$ expansion of relativistic hydrodynamics, where Bjorken flow has been shown to emerge at the leading order. This connection has also been extended to $3+1$ dimensional Gubser like flows in [25, 26]. It remains to be seen whether this Carrollian picture can be connected to an effective parameter dialing as done in the current work.
@@ Acknowledgements
The authors would like to thank Arjun Bagchi, Arpan Das and Kedar Kolekar for useful related discussions. ABan is supported in part by an OPERA grant and a seed grant NFSG/PIL/2023/P3816 from BITS Pilani. He also acknowledges financial support from the Asia Pacific Center for Theoretical Physics (APCTP) via an Associate Fellowship. AG would like to thank BITS Pilani, Pilani Campus for kind hospitality during which a part of this work was done.

### v2

@@ 1 Introduction
Relativistic hydrodynamics has been widely successful in describing collective phenomena in high energy QCD and early universe physics. Based on experimental data obtained from relativistic Heavy Ion Collisions (HICs) conducted both at the RHIC and LHC, we can confirm that a hot and dense nuclear matter is generated during the initial stages of these collisions, commonly referred to as Quark-Gluon Plasma (QGP). It is astonishing that many physical phenomena associated to QGP is well approximated by the almost perfect, strongly-coupled fluid picture of high energy particle processes in these regimes. The problem of solving hydrodynamics in $(1+1)$ dimensions has subsequently receieved particular attention since for very high energy particles, rotational invariance is approximately preserved along the transverse directions [^3] However, rotational symmetry, while keeping the motion independent of azimuthal coordinates, can still allow radial expansions, making the actual system $(1+2)$ dimensional. . Depending on the rapidity dependence of such a flow, one could classify the associated heavy-ion collision processes.
In general, there are two broad classes of solutions deemed most important in this regime, namely the ones with and without manifest boost invariance. The boost invariant regime for a perfect fluid was developed by Bjorken [1] (see also [2]) in his seminal work, which applies beautifully to the central rapidity region of high energy collisions of relativistic heavy ions. Due to the symmetry protected nature, one can easily determine the four-velocity profile for the beam in question. Although elegant in formulation, the Bjorken boost invariant predictions are hardly ever important for experiments. In such cases, still considering a highly idealised situation, Gaussian rapidity distributions are thought to dominate the data for the dense hot matter produced in high energy collisions, giving rise to what is know as the Landau flow [3]. This has been experimentally observed in high energy scattering events [4, 5, 6, 7, 8, 9] [^4] More extensive discussions on Landau flows appear in a number of works, including [10, 11, 12, 13, 14, 15, 16]. This list is by no means exhaustive and the interested reader is requested to look at the references of these.. However lucrative, exact analytic solutions of fluid equations in $(1+1)$ dimensions remain scarce, and general solutions spanning all parameter space is still hard to find even for benchmarking purposes.
@@ 2 The set-up: Hydrodynamic equations
With the light-cone coordinates the partial derivatives become, $\partial_{\pm}=\frac{\partial}{\partial z_{\pm}}\equiv\frac{1}{2}\left(\frac{\partial}{\partial t}\pm\frac{\partial}{\partial z}\right)$ and the hydrodynamic equation takes the form,
$\partial_{\pm}T^{01}+\frac{1}{2}\partial_{+}(T^{11}\pm T^{00})-\frac{1}{2}\partial_{-}(T^{11}\mp T^{00})=0.$ (4)
@@
$y\equiv\frac{1}{2}\ln\left(\frac{\epsilon+p_{z}}{\epsilon-p_{z}}\right)=\frac{1}{2}\ln\left(\frac{1+v}{1-v}\right),$ (5)
where $p_{z}$ is the momentum along $z$ direction (direction of the flow). Substituting the definition of velocity and (2) in (4) we get,
@@ 3 Interpolating flows: Boost and beyond
$y=\frac{1}{2}\ln\left(\frac{\epsilon+p_{z}}{\epsilon-p_{z}}\right)=\frac{1}{2}\ln\left(\frac{t+z}{t-z}\right).$ (9)
@@
which re-confirms that in the $A\rightarrow 0$ limit we do get back the linear term  [^7] Interested readers can check that the above is the expansion one gets if one expands the exact solution.. Another point to be made here is about the dimensionality of $A$. Since rapidity $y$ is a dimensionless parameter, eq(12) tells us that we have freedom to choose the dimensions of the function ff. [^8] If we are being stringent then in (17) we need to add another term to make it dimensionally correct. That is, $z=\int(\sqrt{A^{2}\ln(f/f_{0})+1})^{-1/2}df$. Or one can also just work with a dimensionless $f$ to begin with. Consequently the dimensions of $A$ from (13) is: $\left[A\right]=\left[f\right]\left[L^{-1}\right]$. Since for Bjorken flow the ansatz is simply $f_{\pm}=z_{\pm}$, a natural choice is to set $f$ to be of the dimensions of length which makes our interpolation parameter to be dimensionless; which is also intuitive to some extent since the $A$ gives us an effective dial. All in all, we now have a one-parameter family of hydrodynamic solutions that should interpolate, according to the value of $A$, between boost-invariant and non-invariant regimes.
@@
where, $l_{\pm}=\sqrt{A^{2}\ln{f_{\pm}}+1}$. Here, we have tried to keep the form of the expressions similar to the original work [17]; but, note that the definition of $l_{\pm}$ is different here, and so is the extra $\frac{g+1}{g}$ in the expression of pressure. But more importantly, these expressions can be seamlessly interpolated to the boost invariant regime. To round out this discussion, we show how the general solution based on (28) is different from the boost invariant regime, via plotting the value of $y/\eta$ for varying proper time and varying $A$ in Fig. 1 and 2.
@@
With the explicit form of the function $f_{\pm}$ in (28) we can now also write down the explicit velocity profile of the interpolating flow; but, since the form of $f_{\pm}$ isn’t pretty, consequently the functional form of the velocity profile is not very enlightening and we refrain from mentioning it here.
Figure 2: The ratio of $y/\eta$ has been plotted against the proper time $\tau$ for a constant $\eta$ and varying interpolation parameter $A$. Some we see that for smaller and smaller AA values the graph becomes flatter around 1 which is what we expect since we move towards the Bjorken flow limit.
@@ 4 Entropy flow
$\frac{dp}{d\epsilon}=\frac{s}{T}\frac{dT}{ds}=c_{s}^{2}\equiv\text{const.}$ (31)
@@
$\partial_{\pm}\Phi(z_{+},z_{-})\equiv u^{\mp}T=T_{0}e^{-\theta\mp y}.$ (36)
@@
$=\frac{s_{0}}{\sqrt{2g}}\exp{\left[\frac{g-1}{2A^{2}}\left(\left(\frac{1}{g}+A^{2}\theta\right)+\left(\left(\frac{1}{g}+A^{2}\theta\right)^{2}-\frac{A^{4}y^{2}}{g}\right)^{1/2}-4\right)\right]}$
@@
Note that this is in general a gaussian Landau-like distribution in rapidity. The evolution of this entropy flow for different times and different $A$’s have been shown in Fig. 3 and Fig. 4 respectively.
@@
Figure 3: Plot of the entropy flow distribution for a non-boost invariant flow in (41) at different time instances, keeping other parameters fixed. The redder curves are the plots for later time. We see that the distribution flattens as the fluid flow evolves.
Figure 4: Plot of the entropy flow distribution in (41) for a non-boost invariant flow for different value of $A$, keeping other parameters fixed. The redder curves are the plots for smaller AA value. We see that the distribution flattens for small values of $A$, as expected.
@@ 5 General harmonic flow solutions
As discussed in preceding sections, the solution (26) and (28) are one parameter solutions for the equation (7). But we can actually do better. The anstaz (12) solves the equation (7) but a homogeneous version thereof,
@@
One can now reproduce the solutions presented in [17] and [20] by setting $A=1$ and C=0C=0.151515We emphasize this once again that with $C=0$ one cannot go to the Bjorken regime in any limit, not at least just by dialing parameters.
@@ 6 Discussions and conclusions
In this work, we discussed a refinement of the generalised fluid ansatz that unifies the Landau and Bjorken approaches to studying relativistic hydrodynamics. The full equations of motion can be solved for longitudinal flows effectively in $(1+1)$ dimensions, including free parameters which act as dials between different regimes of physical flows. We found out explicit expressions with arbitrary speed of sound for the physical observables in this case, and also worked out the entropy distribution in rapidity evaluated at different freeze-out temperatures. The harmonic property of the rapidity in hydrodynamic regime gives a large parameter space for viable solutions in this case.
This note was merely to clarify the construction of the interpolating ansatz between boost invariant and non-invariant regimes. There are various questions and intricacies that remain in the discussion of exact solutions in $(1+1)$ hydrodynamics. An immediate question, for example, is the nature of the free parameter $A$. The equations (8), (14), and (26) do not directly associate $A$ with a specific physical quantity. However, the plots in figures (2), and (4) where physical quantities vary with $A$ provide valuable insight into its physical interpretation. In principle this parameter measures the distance between the space-time rapidity and the kinematical rapidity, which coincide in the case of boost invariance. As $A$ increases, the flow increasingly deviates from the Bjorken regime, indicating a breakdown of boost invariance. This could also be thought in terms of the rapidity width of the beam in question, since this quantity diverges at the Bjorken limit, which is uniform in rapidity. The more we add finite width corrections, the more we break boost invariance. In this sense $A$ could be thought of as a quantity related (perhaps invrsely) to rapidity widths as well, however one needs to investigate this in a better way.
A very interesting recent geometric perspective into the problem of Bjorken physics comes from Carrollian symmetries [22], which are basically speed of light going to zero limit of the relativistic cousin thereof [23, 24]. The highly ultra-relativistic nature of heavy-ion flows can be approximated as an effective $c\to 0$ expansion of relativistic hydrodynamics, where Bjorken flow has been shown to emerge at the leading order. In their derivation, the authors implement specific gauge choices for their frame fields and connections, that could be speculated to play a role analogous to the interpolation introduced by our parameter $A$. This connection suggests that $A$ may encode an effective velocity parameter, related to the geometric structures, thereby offering a richer interpretation beyond simple interpolation. This effective emergence of boost invariant flows from Carroll manifolds have also been extended to $3+1$ dimensional Gubser flows in [25, 26]. However, it is beyond the scope of the current work to connect this two disparate situations, both of which lead to Bjorken scalings in the limit.
@@ Acknowledgements
The authors would like to thank Arjun Bagchi, Arpan Das and Kedar Kolekar for useful related discussions. ABan is supported in part by an OPERA grant and a seed grant NFSG/PIL/2023/P3816 from BITS Pilani, and further an early career research grant ANRF/ECRG/2024/002604/PMS from ANRF India. He also acknowledges financial support from the Asia Pacific Center for Theoretical Physics (APCTP) via an Associate Fellowship. AG would like to thank BITS Pilani, Pilani Campus for kind hospitality during which a part of this work was done.